# Mediation analysis in longitudinal intervention studies with an ordinal treatment-dependent confounder

### v1

@@ Abstract
In interventional health studies, causal mediation analysis can be employed to investigate mechanisms through which the intervention affects the targeted health outcome. Identifying direct and indirect (i.e. mediated) effects from empirical data, however, becomes complicated if the mediator-outcome association is confounded by a variable itself affected by the treatment. Here, we investigate identification of mediational effects under such post-treatment confounding in a setting with a longitudinal mediator, time-to-event outcome and a trichotomous ordinal treatment-dependent confounder. We show that if the intervention always affects the treatment-dependent confounder only in one direction (monotonicity), the mediational effects are identified up to a sensitivity parameter and derive their empirical non-parametric expressions. The monotonicity assumption can be assessed from empirical data, based on restrictions on the conditional distribution of the treatment-dependent confounder. We avoid pitfalls related to post-treatment conditioning by treating the mediator as a functional entity and defining the time-to-event outcome as a restricted disease-free time. In an empirical analysis, we use data from the Finnish Diabetes Prevention Study to assess the extent to which the effect of a lifestyle intervention on avoiding type 2 diabetes is mediated through weight reduction in a high-risk population, with other health-related changes acting as treatment-dependent confounders.
@@ 1 Introduction
Lifestyle choices play an important role in the prevention of type 2 diabetes (T2D). As T2D causes major health and economic burdens globally, lifestyle interventions to reduce its incidence in high-risk populations is an area of active research (Tuomilehto et al., 2023; Uusitupa et al., 2019; Barry et al., 2017; Haw et al., 2017). Studies of such interventions typically follow a cohort of individuals over time and aim at collecting information on biomarkers and health status evaluations at several follow-up visits (Kivipelto et al., 2013; Knowler et al., 2002; Gunderson et al., 2007). In countries with extensive health registers, data gathered at follow-up visits can be augmented with time-to-event outcomes retrieved from the registers.
@@
Intervention studies are often based on experimental designs where study participants are assigned to treatment and control groups and the target of inference is the causal effect of an intervention on a specific health response. A more elaborate question regards the extent to which the treatment effect might be mediated through an intermediate variable. For example, potential mediating mechanisms of liraglutide treatment on cardiovascular and chronic kidney disease through changes in clinical biomarkers have been previously investigated in a people with diabetes (Buse et al., 2020; Mann et al., 2021).
@@
Questions regarding mediating mechanisms can be addressed within the causal mediation analysis framework, where the total effect of treatment is decomposed into direct and indirect (i.e. mediated) effects (Robins and Greenland, 1992; Pearl, 2001). The estimands of these causal effects can be defined in terms of counterfactuals, i.e. expected outcomes in hypothetical scenarios where the treatment and the mediator are intervened upon to set them at certain values (Rubin, 1974). Under non-trivial conditional independence assumptions, the estimands can be non-parametrically identified from empirical data (Pearl, 2014). Studies using a lifestyle intervention as the treatment and a clinical risk factor as the mediator, however, are especially prone to failing the identifiability assumptions, because the intervention usually aims at inducing behavioural changes which affect the response not only via the intended mediator, but also through other mechanisms. The behavioural change then acts as a treatment-dependent confounder, i.e. a variable that confounds the mediator–outcome relationship and is itself affected by the treatment. In such situations the standard independence assumptions do not suffice to identify natural mediational effects non-parametrically (Avin et al., 2005).
@@
Tchetgen Tchetgen and VanderWeele (2014) showed that identification can be retained under treatment-dependent confounding by further assumptions such as monotonicity of the treatment effect on a binary treatment-dependent confounder. Monotonicity here means that the treatment may have only a positive (or negative) effect on the treatment-dependent confounder. Other assumptions retaining identifiability include independence of counterfactuals of the treatment-dependent confounder or absence of an additive interaction between the mediator and the treatment-dependent confounder (Robins and Richardson, 2010; Tchetgen Tchetgen and VanderWeele, 2014).
@@
Further methodological issues arise when study participants are followed over time and the mediator is a longitudinal process. While causal mediation analysis has been extended to scenarios with longitudinal mediators (Lin et al., 2017; VanderWeele and Tchetgen Tchetgen, 2017; Zheng and van der Laan, 2017), most literature treats the mediator as a vector-valued entity containing the successive measurements of the mediating variable. Such an approach results in high dimensionality if the number of repeated measurements per individual is large, as well as in problems in handling uneven measurement intervals and missing values. Some authors have instead applied functional regression to represent the mediator trajectory as a function (Lindquist, 2012; Coffmann et al., 2023; Zeng et al., 2023). Treating the mediator as a functional entity avoids the problem of dimensionality and allows a flexible usage of its full history at the time the response is evaluated. The repeated measurements in the longitudinal setting can also be used to extract information about individual-level latent properties. For example, Zheng and Liu (2021) relaxed the assumption of no unmeasured mediator–outcome confounding under a longitudinal mediator and a time-to-event outcome by employing a joint modelling framework to estimate and control a common random effect between the two.
@@
Time-to-event outcomes present additional challenges in the causal inference framework (Didelez, 2019). A particular issue arises if causal estimands are defined via measures that condition on previous survival, such as the hazard function. If there exist latent variables affecting survival, conditioning opens a backdoor path from the treatment to future survival through the latent variables. This problem, however, may readily be dealt with by defining the response as an unconditional measure, such as the restricted mean survival time (RMST), i.e. the mean event-free time within a preset time period (Zeng et al., 2023; Royston and Parmar, 2013). If a significant portion of individuals do not experience the event during the study, RMST has the additional benefit over mean survival time that the considered time period can be chosen to make it robust to misspecification of the shape of the unobserved tail of the event-time distribution. In addition, if it is considered possible to have a zero risk for the event, the true mean survival time would be infinite, whereas RMST would remain constrained to the clinically relevant time period chosen.
@@
In this study, we address identification of mediational causal effects in interventional studies with treatment-dependent confounding. We extend the previously presented monotonicity assumption (Tchetgen Tchetgen and VanderWeele, 2014) to a trichotomous treatment-dependent confounder and show that this leads to expressions that can be identified up to a scalar sensitivity parameter. As an empirical application, we consider the effect of an intensive lifestyle intervention on T2D incidence among a high-risk population, based on the Finnish Diabetes Prevention Study (DPS) (Lindström et al., 2003). The aim of the empirical analysis is to quantify the extent to which the effect of the intervention on T2D-free time is mediated through it inducing weight loss. We apply functional regression analysis to represent the body mass index (BMI) trajectory as a functional entity, and adopt a joint modelling framework to control potential latent confounding between the BMI trajectory and T2D incidence.
@@
The structure of this paper is as follows. In section 2, we present the empirical problem motivating this study. In section 3, we define the targeted causal estimands and provide conditions for their identification and the methods we propose to use for their estimation. Specifically, in subsection 3.4, we derive the expressions for the targeted mediational effects with a latent functional mediator, a time-to-event outcome and a trichotomous treatment-dependent confounder. In sections 4 and 5, we present the results of the empirical application and discuss the results along with further considerations of the used methodology and its possible limitations.
@@ 2 Data sources and motivation
The aim of the Finnish Diabetes Prevention Study (DPS) is to assess the effectiveness of intensive lifestyle intervention in preventing and delaying T2D onset in a high-risk population (Eriksson et al., 1999; Tuomilehto et al., 2001; Lindström et al., 2003). The study cohort was enrolled in 1993–1998 and consisted originally of 522 individuals satisfying the eligibility criteria of being overweight ($\hbox{BMI}>25$), aged 40–64 years and classified at the screening visit as having developed impaired glucose tolerance. They were randomly allocated to the intervention and control groups. The active intervention lasted for a maximum of six years (median 4 years) and entailed frequent personalised nutritional counselling and encouragement for physical activity, received mainly at face-to-face counselling sessions. The active intervention was ended by the end of year 2001. The control group was given routine, non-personalised healthy lifestyle advice during the study visits. Post-intervention follow-up visits in the study cohort continued up until 2013, resulting in a median of ten clinical study visits per person. The study participants were tested for T2D in the clinical study visits based on the World Health Organization 1985 criteria for a 2-hour oral glucose tolerance test (World Health Organization, 1985); for a diagnosis, two OGTTs above the T2D threshold were required. In addition, the Finnish Registers for Drug Reimbursements and Drug Purchases were searched for T2D-related drug purchases, extending the time-to-event follow up until the end of year 2018.
@@
A previous analysis of the DPS data showed a 40% lower hazard for T2D and greater weight loss in the intervention group during the first 13 years of follow-up (Lindström, 2013). The intervention group was also observed to increase the healthiness of their lifestyle compared with the controls, being more likely to make favourable changes to their dietary intakes (Lindström, 2013). Moreover, previous analyses have found associations between the lifestyle intervention, physical activity, nutritional components and diabetes incidence in the DPS cohort (Laaksonen et al., 2005; Lindström et al., 2006).
@@
Since obesity is commonly identified as one of the most prominent risk factors for T2D, it is of interest to assess the extent to which the success of the lifestyle intervention can be attributed to its ability to reduce body weight in people with overweight or obesity. Previous literature based on lifestyle intervention studies has argued in favour of a so-called legacy effect, wherein the intervention effects on T2D incidence persist long after the intervention is discontinued and the obtained group differences in risk factors have attenuated. As summarised in (Wilding, 2014), such results have been reported in the major T2D prevention trials such as the DPS (Lindström, 2013), the study conducted in China (Li et al., 2008) and the Diabetes Prevention Program study in the USA (Diabetes Prevention Program Research Group, 2009). These findings motivate us to consider a parameterisation in which the change in the BMI during the early phase of the intervention as the effective mediator. Moreover, we here apply a three-year time window because the differences in BMI between the two groups was mainly obtained during that time window. We will use the restricted survival time as the outcome measure, choosing the maximum time as 15 years to reflect a clinically relevant time-period. The outcome is thus interpreted as the amount of healthy (i.e. T2D-free) years over the first 15 years after the start of the intervention.
@@
In addition to weight reduction, the aim of the DPS intervention was to reach at least moderate physical activity and healthy nutritional composition measured by the intakes of total fats, saturated fats and fibres (Eriksson et al., 1999). These lifestyle factors can be assumed to influence the study participants’ BMI trajectories and also T2D incidence through mechanisms other than weight loss, thus rendering them potential treatment-dependent confounders. We formed a summary variable based on total physical activity and the components of dietary intake to represent the individual’s lifestyle choices influenced by the intervention. The amount of total physical activity was measured at every study visit by self-reporting and the components of dietary intake by three-day food diaries prior to the study visits for the first three (in addition to the baseline) study visits. All the variables from each study visit were standardised with respect to their baseline means and standard deviations and the lifestyle score was computed as the mean over the standardised variables over the three post-baseline study visits. The lifestyle score was then categorised to three levels based on observed tertiles of the calculated baseline variable.
@@
Conducting causal mediation analysis requires controlling for any factors confounding the associations between the treatment, mediator and outcome. We considered age, sex, smoking status and the baseline lifestyle score as potential confounding baseline variables. Age at baseline was categorised to $<45$, 45 to 59 and $\geq 60$ years, and the smoking status was categorised as ‘never’, ‘former’ or ‘current’.
@@ 3 Methods
In this section we present the proposed methodological framework. Sections 3.1 and 3.2 outline the assumed causal model and the estimation targets. In Section 3.3 we give assumptions sufficient to identify the causal estimands from empirical data and the resulting expressions for the direct and indirect effects. In Sections 3.4 and 3.5 we define the parametric models and the strategy for their estimation.
@@ 3.1 Causal estimands
If there is interaction between the treatment and the mediator, the interpretation of the mediational effects depends on the choice of the levels at which the fixed variables are held in each case. The definitions given above lead to interpreting the direct effect as the pure direct effect and the indirect effect as the total indirect effect, as discussed by VanderWeele (2013, 2014). For ease of notation, we hereafter denote the mediator trajectory and its realisation often simply as $M$ and $m$, respectively.
@@ 3.2 Causal model
The causal model is represented graphically by the directed acyclic graph (DAG) shown in Figure 1, describing the assumed causal structure of the relevant variables involved. The main interest lies in the interplay between the treatment $A$ (lifestyle intervention), the mediator trajectory $M$ (BMI trajectory) and the restricted survival time $T$ as the outcome. The paths from the treatment to the survival time bypassing the mediator reflect the direct effect of the treatment on the outcome, while the indirect effect is reflected in the paths traversing from the treatment to the survival time through the mediator.
We assume that there are individual-specific random effects ($R_{0}$,$\boldsymbol{R}$) that affect the outcome through the underlying mediator trajectory. In addition, $R_{0}$ is allowed to have a separate, direct effect, constituting as a frailty for the time-to-event outcome. The treatment-dependent confounder $U$ represents lifestyle behaviour that is influenced by the intervention but is not of primary interest. The variable set $W$ consists of the confounding baseline variables that are considered to affect the mediator trajectory, the restricted survival time and the lifestyle behaviour.
@@
Figure 1: Directed acyclic graph describing the assumed causal mechanism. The effect of treatment $A$ on the restricted survival time $T$ is assumed to be mediated through the underlying mediator trajectory $M(t)$ of which instances at certain time points are observed. The random effect $R_{0}$ is shared between the mediator and the time-to-event outcome, while the random effects $\boldsymbol{R}$ affect only the mediator. The dashed line between $R_{0}$ and $\boldsymbol{R}$ implies that they are correlated. $U$ is the lifestyle behaviour influenced by the intervention and $\boldsymbol{W}$ consists of the baseline confounders. The dashed nodes refer to latent (unobserved) variables.

### v2

@@ Abstract
In interventional health studies, causal mediation analysis can be employed to investigate mechanisms through which the intervention affects the targeted health outcome. Identifying direct and indirect (i.e. mediated) effects from empirical data become complicated, however, when the mediator-outcome association is confounded by a variable itself affected by the treatment. Here, we investigate identification of mediational effects under such post-treatment confounding in a setting with a longitudinal mediator, time-to-event outcome and a trichotomous ordinal treatment-dependent confounder. If the intervention always affects the treatment-dependent confounder only in one direction (monotonicity), we show that the mediational effects are identified up to a stratum-specific sensitivity parameter and derive their empirical non-parametric expressions. The feasibility of the monotonicity assumption can be assessed using empirical data, based on restrictions on the marginal distributions of counterfactuals of the treatment-dependent confounder. We avoid pitfalls related to post-treatment conditioning by treating the mediator as a functional entity and defining the time-to-event outcome as a restricted disease-free time. In an empirical analysis, we use data from the Finnish Diabetes Prevention Study to assess the extent to which the effect of a lifestyle intervention on avoiding type 2 diabetes is mediated through weight reduction in a high-risk population, with other health-related changes acting as treatment-dependent confounders.
@@ 1 Introduction
Lifestyle choices play an important role in the prevention of type 2 diabetes (T2D). As T2D causes major health and economic burdens globally, lifestyle interventions to reduce its incidence in high-risk populations are an area of active research (Tuomilehto et al., 2023; Uusitupa et al., 2019; Barry et al., 2017; Haw et al., 2017). Studies of such interventions typically follow a cohort of individuals over time and aim at collecting information on biomarkers and health status evaluations at several follow-up visits (Kivipelto et al., 2013; Knowler et al., 2002; Gunderson et al., 2007). In countries with extensive health registers, data gathered at follow-up visits can be augmented with time-to-event outcomes retrieved from the registers.
@@
Intervention studies are often based on experimental designs where study participants are assigned to treatment and control groups and the target of inference is the causal effect of an intervention on a specific health response. A more elaborate question involves understanding the extent to which the treatment effect is mediated through an intermediate variable. For example, potential mediating mechanisms of liraglutide treatment on cardiovascular and chronic kidney disease through changes in clinical biomarkers have been previously investigated in people with diabetes (Buse et al., 2020; Mann et al., 2021).
@@
Mediating mechanisms can be addressed within the causal mediation analysis framework, where the total effect of treatment is decomposed into direct and indirect (i.e. mediated) effects (Robins and Greenland, 1992; Pearl, 2001). The estimands of these causal effects can be defined in terms of counterfactuals, i.e. expected outcomes in hypothetical scenarios where the treatment and the mediator are intervened upon to set them at certain values (Rubin, 1974). Under non-trivial conditional independence assumptions, the estimands can be non-parametrically identified from empirical data (Pearl, 2014). Studies using a lifestyle intervention as the treatment and a clinical risk factor as the mediator, however, are especially prone to failing the identifiability assumptions, because the intervention often induces behavioural changes that affect the response through both the intended mediator and other mechanisms. The behavioural change then acts as a treatment-dependent confounder, i.e. a variable that confounds the mediator–outcome relationship and is itself affected by the treatment. In such situations the standard independence assumptions do not suffice to identify natural mediational effects non-parametrically (Avin et al., 2005).
@@
Tchetgen Tchetgen and VanderWeele (2014) showed that identification can be retained even under treatment-dependent confounding by further assumptions such as monotonicity of the treatment effect on a binary treatment-dependent confounder. Monotonicity here means that the treatment may have only a positive (or negative) effect on the treatment-dependent confounder. Other assumptions retaining point identification include independence of counterfactuals of the treatment-dependent confounder or absence of an additive interaction between the mediator and the treatment-dependent confounder (Robins and Richardson, 2010; Tchetgen Tchetgen and VanderWeele, 2014). Without any additional assumptions, the mediational effects under an observed discrete treatment-dependent confounder are still partially identifiable. This means that lower and upper bounds for the estimates can be obtained (Miles et al., 2017).
@@
Further methodological issues arise when study participants are followed over time and the mediator is a longitudinal process. While causal mediation analysis has been extended to scenarios with longitudinal mediators (Lin et al., 2017; VanderWeele and Tchetgen Tchetgen, 2017; Zheng and van der Laan, 2017), most literature treats the mediator as a vector-valued entity containing successive measurements of the mediating variable. Such an approach results in high dimensionality if the number of repeated measurements per individual is large and also poses challenges in handling uneven measurement intervals and missing values. Some authors have instead applied functional regression to represent the mediator trajectory as a function (Lindquist, 2012; Coffmann et al., 2023; Zeng et al., 2023). Treating the mediator as a functional entity avoids the problem of dimensionality and enables flexible use of its full history at the time the response is evaluated. The repeated measurements in the longitudinal setting can also be used to extract information about individual-level latent properties. For example, Zheng and Liu (2021) relaxed the assumption of no unmeasured mediator–outcome confounding under a longitudinal mediator and a time-to-event outcome by employing a joint modelling framework to estimate and control a common random effect reflecting an unobserved confounder between the two.
@@
Time-to-event outcomes pose additional challenges in the causal inference framework (Didelez, 2019). A particular issue arises when causal estimands are defined by measures that condition on prior survival, such as the hazard function. If there exist latent variables affecting survival, conditioning opens a backdoor path from the treatment to future survival through the latent variables. This issue can be addressed by defining the response as an unconditional measure, such as the restricted mean survival time (RMST), i.e. the mean event-free time within a preset time period (Zeng et al., 2023; Royston and Parmar, 2013). If a significant portion of individuals do not experience the event during the study, RMST has the additional benefit over the mean survival time of allowing the time period to be chosen so that it remains robust to misspecification of the unobserved tail of the event-time distribution. In addition, if it is considered possible to have a zero risk for the event, the true mean survival time would be infinite, whereas RMST would remain constrained to the chosen clinically relevant time period.
@@
In this study, we address identification of mediational causal effects in interventional studies with treatment-dependent confounding. We extend the previously presented monotonicity assumption (Tchetgen Tchetgen and VanderWeele, 2014) to a trichotomous treatment-dependent confounder and show that this results in expressions identifiable up to a stratum-specific sensitivity parameter. Our approach is similar to the partial identification in Miles et al. (2017) but imposes restrictions on the unobserved joint distribution of the counterfactuals of the treatment-dependent confounder, leading to a necessary condition for their marginals. As the marginals can be estimated from observed data, the feasibility of the monotonicity assumption can be empirically assessed. As an application, we consider the effect of an intensive lifestyle intervention on T2D incidence among a high-risk population, based on the Finnish Diabetes Prevention Study (DPS) (Lindström et al., 2003). The aim of the empirical analysis is to quantify the extent to which the effect of the intervention on T2D-free time is mediated through it inducing weight loss. We apply functional regression analysis to represent the body mass index (BMI) trajectory as a functional entity and use a joint modelling framework to control potential latent confounding between the BMI trajectory and T2D incidence.
@@
The paper is structured is as follows. Section 2, presents the empirical problem motivating this study. Section 3, defines the targeted causal estimands and provides conditions for their identification and the methods we propose to use for their estimation. Sections 4 and 5, present the results of the empirical application and discuss the results along with further considerations of the used methodology and its possible limitations.
@@ 2 Data sources and motivation
The aim of the Finnish Diabetes Prevention Study (DPS) is to assess the effectiveness of an intensive lifestyle intervention in preventing and delaying T2D onset in a high-risk population (Eriksson et al., 1999; Tuomilehto et al., 2001; Lindström et al., 2003). The study cohort was enrolled between 1993 and 1998 and originally consisted of 522 individuals. The eligibility criteria required being overweight ($\hbox{BMI}>25$), aged 40–64 years and having impaired glucose tolerance at the screening visit. The study participants were randomly allocated to intervention and control groups. The active intervention lasted a maximum of six years (median 4 years), involving frequent personalised nutritional counselling and encouragement for physical activity, primarily through face-to-face sessions. The active intervention ended in 2001. The control group was given routine, non-personalised healthy lifestyle advice during the study visits. Post-intervention follow-up visits continued until 2013, with a median of ten clinical study visits per person. The participants were tested for T2D in the clinical study visits using the World Health Organization’s 1985 criteria for a 2-hour oral glucose tolerance test (OGTT) (World Health Organization, 1985), with a diagnosis requiring two OGTTs above the threshold. In addition, the Finnish Registers for Drug Reimbursements and Drug Purchases were searched for T2D-related drug purchases, extending the time-to-event follow up until the end of 2018.
@@
A previous analysis of the DPS data showed a 40% lower hazard for T2D and greater weight loss in the intervention group during the first 13 years of follow-up, and also showed greater improvements in their lifestyle compared with the control group, particularly in dietary intakes (Lindström et al., 2013). Moreover, previous analyses have found associations between the lifestyle intervention, physical activity, nutritional components and diabetes incidence in the DPS cohort (Laaksonen et al., 2005; Lindström et al., 2006).
@@
Since obesity is one of the most prominent risk factors for T2D, it is of interest to assess the extent to which the success of the lifestyle intervention can be attributed to its ability to reduce body weight in people with overweight or obesity. Previous literature on lifestyle intervention studies has suggested a so-called legacy effect, wherein the intervention’s impact on T2D incidence persist long after the intervention ends and the obtained group differences in risk factors have diminished. As summarised by Wilding (2014), such results have been reported in the major T2D prevention trials, including the DPS (Lindström et al., 2013), the study conducted in China (Li et al., 2008) and the Diabetes Prevention Program in the USA (Diabetes Prevention Program Research Group, 2009). These findings motivated us to consider the change in the BMI during the early phase of the intervention as the effective mediator. Moreover, we apply a three-year time window because the majority of differences in BMI between the two groups occurred during this period. We will use the restricted survival time as the outcome measure, choosing the maximum time as 15 years, reflecting a clinically relevant time period. The outcome is thus interpreted as the number of healthy (i.e. T2D-free) years during the first 15 years after intervention onset.
@@
In addition to weight reduction, the DPS intervention aimed at moderate physical activity and healthy nutritional composition measured by intakes of total fats, saturated fats and fibre (Eriksson et al., 1999). These lifestyle factors can be assumed to influence the study participants’ BMI trajectories and also T2D incidence through mechanisms other than weight loss, thus rendering them potential treatment-dependent confounders. We created a summary variable to represent individuals’ lifestyle choices influenced by the intervention, combining total physical activity and the dietary intake components. The amount of total physical activity was measured by self-reports at every study visit and the components of dietary intake by three-day food diaries prior to the study visits for the first three (in addition to the baseline) study visits. All variables from each study visit were standardised with respect to their baseline means and standard deviations and the lifestyle score was computed as the mean over the standardised variables over the three post-baseline study visits. The lifestyle score was then categorised to three levels based on the observed tertiles of the score at baseline.
@@
Causal mediation analysis requires controlling for any factors confounding the relationships between the treatment, mediator and outcome. We considered age, sex, smoking status and the baseline lifestyle score as potential confounding baseline variables. Age at baseline was categorised as $<45$, 45 to 59 and $\geq 60$ years, while the smoking status was categorised as ‘never’, ‘former’ or ‘current’.
@@ 3 Methods
In this section we present the proposed methodological framework. Sections 3.1 and 3.2 outline the assumed causal model and the estimation targets. In Section 3.3 we give assumptions sufficient to identify the causal estimands from empirical data and the resulting expressions for the direct and indirect effects. In Sections 3.4 and 3.5 we define the parametric models and describe the strategy for their estimation.
@@ 3.1 Causal estimands
If there is interaction between the treatment and the mediator, the interpretation of mediational effects depends on the choice of the levels at which the fixed variables are held in each case. The definitions (1) lead to interpreting the direct effect as the pure direct effect and the indirect effect as the total indirect effect, as discussed by VanderWeele (2013, 2014). For ease of notation, we hereafter denote the mediator trajectory and its realisation often simply as $M$ and $m$, respectively.
@@ 3.2 Causal model
The causal model is represented graphically by the directed acyclic graph (DAG) shown in Figure 1, describing the assumed causal structure of the relevant variables involved. The main interest lies in the interplay between the treatment $A$ (lifestyle intervention), the mediator trajectory $M$ (BMI trajectory) and the restricted survival time $T$ as the outcome. The paths from the treatment to the survival time bypassing the mediator represent the direct effect of the treatment on the outcome, while the indirect effect follows the paths traversing from the treatment to the survival time through the mediator. The treatment-dependent confounder $U$ represents lifestyle behaviour that is influenced by the intervention but is not of primary interest. The variable set $W$ consists of the confounding baseline variables that are considered to affect the mediator trajectory, the restricted survival time and the lifestyle behaviour.
We assume that there are individual-specific random effects ($R_{0}$,$\boldsymbol{R}$) that affect the outcome through the underlying mediator trajectory. In addition, $R_{0}$ is allowed to have a separate, direct effect, constituting frailty for the time-to-event outcome, and furthermore, a confounder for the causal mechanism. The random effects can be seen as some properties of the individuals which could in principle be measured and adjusted for and, although unobserved in reality, they can be learned through their influence on the repeated measurements of the mediator. A similar approach has previously been adopted to estimate the mediational effects of certain drug treatments on overall survival mediated by CD4 cell count (Zheng and Liu, 2021).
@@
Figure 1: Directed acyclic graph describing the assumed causal mechanism within an individual. The effect of treatment $A$ on the restricted survival time $T$ is assumed to be mediated through the underlying mediator trajectory $M(t)$ of which $n$ repeated measurements $m_{1}^{(o)},\dots,m_{n}^{(o)}$ are observed over time. The random effect $R_{0}$ is shared between the mediator $M(t)$ and the time-to-event outcome $T$, while the random effects $\boldsymbol{R}$ affect only the mediator. The dashed line between $R_{0}$ and $\boldsymbol{R}$ implies that they are correlated. $U$ is the treatment-dependent confounder and $\boldsymbol{W}$ consists of the baseline confounders. The dashed nodes refer to latent (unobserved) variables.